Here is my problem.
The perimeter of the rectangle is 24 and the perimeter of the triangle is 23
Here are the sides of the rectangle:
side 1- x
side 2- 3y
side 3- x
side 4- 3y

here are the sides of the isosoles triangle:
side 1- x+1
side 2- x+y
side 3- y+6
Write a system of linear equations to represent this problem.
How would I do that?

Answer by stanbon(75887)

Here is my problem.
The perimeter of the rectangle is 24 and the perimeter of the triangle is 23
Here are the sides of the rectangle:
side 1- x
side 2- 3y
side 3- x
side 4- 3y
-------------------
Equation:
1-x + 2-3y+3-x + 4-3y = 24
-2x -6y + 10 = 24
-2x - 6y = 14
x + 3y = -7
--------------------
here are the sides of the isosoles triangle:
side 1- x+1
side 2- x+y
side 3- y+6
Equation:
x+1 + x+y + y+6 = 23
2x + 2y + 7 = 23
2x + 2y = 16
x + y = 8
----------------------------
Write a system of linear equations to represent this problem.
x + 3y = -7
x + y = 8
----
Subtract the 2nd equation from the 1st:
2y = -15
y = -15/2
------
Since x + y = 8, substitute and solve for "x":
x -15/2 = 16/2
x = 31/2
-------------------
Cheers,
Stan H.